OpenStudy (anonymous):

Brayden is investing $1,617 in three different stocks: Zycodec, Unix, and Xenon. He will invest 40% of the total between Zycodec and Xenon in the Unix stock. Twice as much as the investment in Zycodec will be given to Xenon. How much will Brayden invest in the Unix stock?

OpenStudy (anonymous):

I broke the question down but i just don't understand how to even begin calculating it! someone help please

OpenStudy (anonymous):

This isn't a question about STOCKS. It's a brain dead algebra question

OpenStudy (anonymous):

That isn't what i asked...

OpenStudy (anonymous):

Someone HELPFUL please respond

OpenStudy (ganpat):

its lil confusing for me..

OpenStudy (anonymous):

I know right! i'll break it down, maybe you will know how to calculate it. He has 1,617 $ to spend there are 3 stocks I have to find the toltal amount he invested in Zycodec & xenon because whatever the total of those two are, he invest 40% of the total on the Unix

OpenStudy (ganpat):

so , x + y + +z = 1617.. 40 % ( x + y ) = z x + y = 100 * z / 40 so (100 * z)/ 40 + z = 1617 140 z = 1617 * 40 so z= 462 $.. this can be possibility..

OpenStudy (stephstar101):

I got $385 for z....

OpenStudy (ganpat):

@stephstar101 : can u plz show steps ?

OpenStudy (anonymous):

Stephstar101, that is one of my answer choices. But how did you calculate it?

OpenStudy (stephstar101):

it's not the answer, because it was the amount for Zycodec, Unix stock should be $462, is that one answer?

OpenStudy (anonymous):

Yes that is one of the answer choices too

OpenStudy (stephstar101):

Okay I'll explain in a sec

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (stephstar101):

So you need to start with 3 different equations based off of what the question asks. The variables assigned correlate with the company name. Equation 1: Z+U+X=1617 <- from the first sentence, he'll invest this much total in 3 stocks Equation 2: .40(Z+X)=U <- comes from sentence 2, 40% of the sum of Z+X will equal U Equation 3: 2Z=X <-2 times Z is the amount given to X Let's start with the math! .40(Z+2Z)=U <- replace the X from equation 2 with the X amount from equation 3 .40(3Z)=U <- add the Zs 1.2Z=U <- multiply .40 and 3 to get 1.2 Z+1.2Z+2Z=1617 <- plug in equation 3 and the 1.2Z you just solved for into equation 1 4.2Z=1617 <-add Zs together Z=$385 <- divide 1617 by 4.2 to get the amount for company Zycodec Plug it into .40(3Z)=U from a previous step. .40(3*385)=U U=$462, and that's your answer!

OpenStudy (ganpat):

@stephstar101 : so 462 is right answer ??

OpenStudy (stephstar101):

Pretty much, since Unix is U, and that's what it's looking for.

OpenStudy (anonymous):

That is a long equation But thank you for explaining... I'm copying it down for my notes!

OpenStudy (stephstar101):

Yeah, there are 3 equations too many >.< But I hope it helped a little going through ALL the steps. Probably the hardest part is understanding the question and finding the equations

OpenStudy (anonymous):

Hahah i know, i read it like 15 times!

OpenStudy (anonymous):

do you by any chance know anything about calculating the amount of a loan after it's deferment period?

OpenStudy (stephstar101):

Ahhhh, finance isn't my strong spot, but ask the question and maybe someone else can also help! I'll try my best to help as well!

OpenStudy (ganpat):

what i tried is simpler one, with reduced equations.. try that one..